# In the 39th Conference on Neural Information Processing Systems (NeurIPS 2025) Workshop: Embodied World Models for Decision Making

### v2

@@ 7 Conclusion
The pursuit of more capable and autonomous agents has led the research community to recognize the critical role of environments in shaping intelligence. The current emphasis on scaling the fidelity and size of these environments, while valuable, overlooks a more crucial dimension: structural complexity. In complex multi-agent domains, the primary obstacle to learning is not merely the sparsity of rewards but the sparsity of coherent, long-horizon tasks. An embodied world model for decision-making must therefore model not just the physics of the world, but also the structure of tasks within it.
@@
This paper has argued that the most effective path forward is to build environments with explicit hierarchical scaffolding. Through a systematic review of the 2024 literature in multi-agent soccer, we have shown a clear trend towards hybrid systems that integrate symbolic planning and hierarchical decomposition with MARL. These methods succeed precisely because they provide the structural priors necessary to make the problem of strategic coordination tractable.
We have extended this position by proposing a new class of structured task environments that treat hierarchical decomposition as a first-class feature, and have argued that Large Language Models can serve as dynamic, generative world models of tasks to create this structure on the fly. This paradigm directly enables research into long-horizon planning, sim-to-real transfer, and language-guided decision making. To scale agents to complex, strategic domains, we must first scale the environments to support complex, strategic learning. Our position is a call to action for the community of environment builders. We must move beyond creating mere physics sandboxes and begin engineering structured learning curricula directly into our benchmarks. Only by scaffolding the world can we expect our agents to build the towering capabilities of goal-directed, embodied intelligence we envision.

### v3

@@ 7 Conclusion
The pursuit of more capable and autonomous agents has led the research community to recognize the critical role of environments in shaping intelligence. The current emphasis on scaling the fidelity and size of these environments, while valuable, overlooks a more crucial dimension: structural complexity. This paper analyzed the limitations of end-to-end MARL in complex, long-horizon domains, identifying task sparsity as a fundamental barrier. An embodied world model for decision-making must therefore model both the physics of the world and the structure of tasks within it.
@@
Our review of the 2024 literature in multi-agent soccer revealed a clear trend towards hybrid systems that integrate symbolic planning and hierarchical decomposition with MARL. These methods succeed precisely because they provide the structural priors necessary to make the problem of strategic coordination tractable.
Building on this analysis, we proposed the Structured Task Environment (HTE) framework, which externalizes task hierarchy into the world model itself. We further proposed a novel implementation using LLMs as dynamic, generative world models of tasks to create this structure on the fly. This paradigm directly enables research into long-horizon planning, compositional generalization, and language-guided decision-making. To scale agents to complex, strategic domains, we must first scale the environments to support complex, strategic learning. Future work will focus on implementing the proposed LLM-based task API and validating the framework’s impact on sample efficiency and generalization in benchmark domains.
Appendix A Technical appendices and supplementary material
To make our proposal concrete, we outline a minimal specification for the Hierarchical Task API described in Section 5.2.
A.1 A.1 Task schema
A task can be defined by a data structure that specifies its components. A planner (like an HTN or LLM) would populate these structures, which the environment would then use to manage the learning process.
class HierarchicalTask:
def init(self, name, parent_task, subtasks):
self.name = name # e.g., "GiveAndGo"
self.parent = parent_task # Task object or None
self.subtasks = subtasks # Ordered list of Task objects
self.status = "inactive" # inactive, active, completed, failed
def check_preconditions(self, world_state):
    # Logic to check if this task can start
    # e.g., for "PassToTeammate", check "agent_has_ball"
    pass
def check_termination(self, world_state):
    # Logic to check if this task is done
    # e.g., for "PassToTeammate", check "teammate_has_ball"
    pass
def get_intrinsic_reward(self, world_state):
    # Return reward if check_termination() is true
    pass
A.2 A.2 Example: HDDL-to-API mapping
As noted in Section 4.2, an HDDL planner can generate the task structure. A high-level HTN method for a (ScoreGoal) task might be:
(:method (ScoreGoal)
:subtasks (and
(AcquireBall ?agent1)
(DribblePastOpponent ?agent1 ?opponent)
(PassToTeammate ?agent1 ?agent2)
(ShootAtGoal ?agent2)
This would map to our API as follows:
Create sub-tasks (as "leaf" nodes)
task_acquire = HierarchicalTask("AcquireBall", ...)
task_dribble = HierarchicalTask("DribblePastOpponent", ...)
task_pass = HierarchicalTask("PassToTeammate", ...)
task_shoot = HierarchicalTask("ShootAtGoal", ...)
Create the parent "composite" task
task_score = HierarchicalTask(
name="ScoreGoal",
parent=None,
subtasks=[task_acquire, task_dribble, task_pass, task_shoot]
A.3 A.3 Standardized soccer task definitions
Here are two standardized task definitions using this structure.
1. Task: GainPossession
  • Description: A single agent must gain control of the ball.
  • name: "GainPossession"
  • subtasks: [ MoveToBall, ControlBall ]
  • preconditions(MoveToBall): ball.is_loose()
  • termination(MoveToBall): $distance(agent,ball)<0.5m$
  • termination(ControlBall): agent.has_ball == True
  • get_intrinsic_reward(): +0.5+0.5 on MoveToBall completion, +1.0+1.0 on ControlBall completion.
2. Task: GiveAndGo (Multi-agent)
  • Description: Agent A passes to Agent B, runs past a defender, and receives a return pass.
  • name: "GiveAndGo"
  • subtasks: [ Pass(A, B), MoveToOpenSpace(A), Pass(B, A) ]
  • preconditions(Pass(A, B)): agent(A).has_ball == True
  • termination(Pass(A, B)): agent(B).has_ball == True
  • preconditions(MoveToOpenSpace(A)): agent(B).has_ball == True (Starts after pass)
  • termination(MoveToOpenSpace(A)): agent(A).is_in_open_space == True
  • get_intrinsic_reward(): +1.0+1.0 for each sub-task completion.